OpenStudy (anonymous):

how would I solve sin (2arctan 12/5) ?

OpenStudy (anonymous):

do I have to use a double angle identity ?? if so how do I do that

OpenStudy (cwrw238):

first use your calculator to find arctan 12/5

OpenStudy (anonymous):

im in a situation where I cant use a calculator actually but I do have all the values for y x and r

OpenStudy (anonymous):

x is 5 y is 12 and r is 13

OpenStudy (cwrw238):

yes thats the 13-12-5 right angled triangle

OpenStudy (anonymous):

mhm so sin is 12/13 but im unsure where to go next

OpenStudy (cwrw238):

sin 2x = 2sinx cosx

OpenStudy (anonymous):

but wouldn't I use tan 2x + 2tanx/ 1-tan squared x ?

OpenStudy (cwrw238):

if sin is 12/13 whats the cos?

OpenStudy (cwrw238):

no

OpenStudy (anonymous):

5/13

OpenStudy (anonymous):

so I wouldn't use the double angle id for tan ? id use it for sin?

OpenStudy (jdoe0001):

\(\bf sin\left[2tan^{-1}\left(\cfrac{12}{5}\right)\right]\implies sin(2\theta)\implies 2sin(\theta)cos(\theta)\\ \quad \\ \quad \\ a= 5\qquad b=12\qquad c=13\qquad thus\\ \quad \\ cos(\theta)=\cfrac{5}{13}\qquad sin(\theta)=\cfrac{12}{13}\qquad 2sin(\theta)cos(\theta)=2\left(\cfrac{12}{13}\right)\left(\cfrac{5}{13}\right)\)

OpenStudy (anonymous):

omg. thank u so much

OpenStudy (cwrw238):

ok so plug into the formula for sin 2x as jdoe0001 has done

OpenStudy (anonymous):

definitely keepin this on hand thanks everyone !!!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

arctan 12/5 means 'the angle whose tangent is 12/5'

OpenStudy (jdoe0001):

yes, in this case \(\bf \theta\)

OpenStudy (jdoe0001):

whatever angle that happen to be